If tan θ=15/8, find the values of all T-ratios of θ.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given function: tan θ=15/8


We know that tan θ = BC/AB = perpendicular/base
= 15/8
By Pythagoras Theorem:
AC^2 = AB^2 + BC^2

= 64k^2 + 225k^2
= 289k^2
AC = 17k
Find other T-rations using their definitions:

Sin𝜃= BC/AC = 15/17

cos𝜃= AB/AC = 8/17
cosec 𝜃= 1/sin𝜃= 17/15
sec𝜃= 1/cos𝜃= 17/8
cot𝜃= 1/tan𝜃= 8/15

Thanks